The tickets for a dance recital cost $5.00 for adults and $2.00 for children.If the total number of tickets sold was 295 and the total amount collected was $1,220, how many adult tickets were sold?
Answer by TimothyLamb(4379)

x = adult tickets sold
y = child tickets sold
---
x + y = 295
5.00x + 2.00y = 1220
---
put the system of linear equations into standard form
---
x + y = 295
5.00x + 2.00y = 1220
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = adult tickets sold = 210
y = child tickets sold = 85
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---